OpenStudy (anonymous):

Can someone help me find the solution to this quadratic equation: 9x2 - 12 + 2 =0 thanks!

OpenStudy (anonymous):

(the 2 after the x is supposed to be squared)

OpenStudy (anonymous):

3^2 x^2=10 square root both sides, we get 3x=square root 10 x=square root 10/3

OpenStudy (anonymous):

\[9x ^{2}-12+2=0,9x ^{2}-10=0,x ^{2}=\frac{ 10 }{9 },x=\pm \frac{ \sqrt{10} }{3 }\]

OpenStudy (anonymous):

thank you both so much!